# The Energy Cost of Artificial Intelligence of Things Lifecycle

### v2

@@ Abstract
Artificial Intelligence (AI) coupled with the existing Internet of Things (IoT) enables more autonomous operations across various economic sectors. While this paradigm shift results in increased energy consumption it is difficult to quantify the end-to-end energy consumption of such systems with the conventional metrics as they either focus on the communication, the computation infrastructure or model development. To address this, we propose a new metric, the Energy Cost of AI lifecycle ($eCAL$). eCAL captures the energy consumption throughout the architectural components and lifecycle of an AI-powered wireless system by analyzing the complexity of data collection and manipulation in individual components and deriving overall and per-bit energy consumption. We show that the better a model and the more it is used, the more energy efficient an inference is. For an example Artificial Intelligence of Things (AIoT) configuration, eCAL for making $100$ inferences is $2.73$ times higher than for $1000$ inferences. Additionally, we developed a modular open source simulation tool to enable researchers, practitioners, and engineers to calculate the end-to-end energy cost with various configurations and across various systems, ensuring adaptability to diverse use cases.
@@
AIoT lifecycle, Energy Consumption, Carbon Footprint, Metric, Methodology
@@ I Introduction
Next-generation networks aim to deliver unprecedented levels of connectivity and distributed intelligence, facilitating seamless interactions between a massive number of devices, services, and applications. The trend is especially noticeable in the IoT where devices are becoming increasingly more intelligent [1], from smart home appliances [2] to industrial machinery [3], smart cities [4], intelligent car [5], and beyond [6, 7]. With the help of Artificial Intelligence (AI), we are moving away from conventional IoT applications towards AIoT systems [8] designed to achieve more streamlined and autonomous operations. For example, in future smart factories, the integration of AI and IoT will enable real-time monitoring and optimization of production processes, predictive maintenance, and enhanced decision-making capabilities.
@@
While AIoT systems enable unprecedented automation and agility, their reliance on AI implies additional energy consumption and increased carbon emissions [9]. Especially in the cases when models with many parameters are employed, such as large language models, these energy and carbon costs may be significant [10]. In the light of climate challenges, the general AI research community’s efforts are intensifying to better estimate the costs of training [11] and using [12] AI, the Carbon Footprint (CF) of LLMs [13] and find ways to scale models efficiently [14]. Furthermore, the networking and IoT communities that are increasingly relying on AI techniques are also investigating pathways towards net-zero carbon emissions [15], analyzing the CF of various learning techniques [16] and data modalities while also developing various approaches to optimizing energy efficiency through hardware-software optimization [17], scheduling [18], more efficient model design [19] and energy and carbon consumption testing [20].
@@
Depending on the scope of the study, various metrics are employed to understand energy efficiency and CF. As well noted in [15], the traditional ITU standardized Energy-per-Bit $[J/b]$ metric ’will no longer be able to reflect the environmental impact of the modern mobile services, especially network AI-enabled smart services’. For instance, [17] considers normalized energy as well as energy for buffering in $[J]$, electricity $[Wh]$ and $[CO_{2}eq]$ for the various phases involved in federated learning (FL) edge systems [15, 16], $[J]$ and GFLOPS of the neural architectures [19] and energy $[Wh]$ for the scheduled loads in [18]. However, these metrics are not specifically designed to capture the energy efficiency of a system or its parts.
Recently, a number of metrics appeared addressing environmental friendliness, such as accuracy per consumption (APC), which captures the performance of an AI model together with its environmental impact [21]. Nevertheless, while the metrics proposed in [21] are powerful in terms of assessing the energy/performance trade-offs of models, they are less suitable in assessing the energy efficiency of a system such as AIoT. To this end, a metric somewhat similar to the Energy-per-Bit $[J/b]$ would be more suitable for measuring energy efficiency of an AI-based communication system, similarly as some other efficiency metrics that have been proposed in wireless in particular and engineering and economy in general [22]. For instance, spectral efficiency, measured in $[b/s/Hz]$ [23] provides means of calculating the amount of data bandwidth available in a given amount of spectrum, and lifecycle emissions for vehicles, measured in $[g/km]$ [24], enables computing the CF in grams per kilometer.
@@
Following these observations, we aimed to find a suitable energy efficiency metric that would be simple and general and would holistically quantify the environmental cost of inference carried out in AIoT systems. The contributions of this paper are as follows:
@@
  • We propose a new metric, the Energy Cost of AIoT Lifecycle (eCAL), measured in $[J/b]$, that captures the overall energy cost of generating an inference in an AIoT system. Unlike capturing the energy required for transmitting bits that is done by the Energy-per-Bit $[J/b]$ metric, or $[J]$ for AI model complexity, the proposed eCAL captures the energy consumed by all the data collection and manipulation components in an AIoT system during the lifecycle of a trained model enabling inference.
@@
  • Following the Open Systems Interconnection (OSI) and Machine Learning Operations (MLOps) [25] architectures, we devise a methodology for determining the eCAL of an AIoT system and develop an open source modular and extensible simulation tool [^1] https://github.com/sensorlab/eCAL, namely, eCAL calculator. The methodology breaks down the AIoT system into different data manipulation components, and, for each component, it analyzes the complexity of data manipulation and derives the overall energy consumption.
@@
  • Using the proposed metric, we demonstrate that the better a model is and the more it is used, the more energy-efficient an inference is. For an example AIoT configuration, the energy consumption per bit for making $100$ inferences is $2.73$ times higher than for $1000$ inferences. This can be explained by the fact that the energy cost of data collection, preprocessing, training, and evaluation, all necessary to develop an AI model are spread across more inferences.
@@
The paper is structured as follows. We first review the relevant literature in Section II. We then describe the lifecycle of an AIoT system, define eCAL and the methodology for deriving it in Section III. Subsequently, the derivation of the energy consumption formulas for different data manipulation components involved in the AIoT lifecycle are provided in Sections IV to VII. In Section VIII, the proposed metric, eCAL, is derived and analyzed, demonstrating its utility in capturing the overall energy cost of an AIoT system over its lifecycle. Finally, Section IX concludes the paper and outlines future research directions.
@@
Proposed Metric eCAL $J/b$ Quantifies the energy consumption of AI inference in AIoT, considering both communication and computation energy costs relative to application data.
@@ II Related Work
The IoT systems increasingly leverage AI techniques to optimize automation and enhance operational efficiency, essentially turning into so called AIoT systems. However, their growing dependence on AI introduces significant energy and carbon costs [9], intensifying environmental concerns [13]. In response, the AI research community is placing greater emphasis on understanding and then formalizing methods to assess the environmental impact of the increasing AI utilization. Due to the complexity of state of the art neural architectures, in many cases the energy cost of training is computed after the training by measuring the performed computation through interfaces such as the performance application programming interface [29]. Furthermore, [11] shows that the energy footprint of inference is more studied than the one for training and highlights below 70% accuracy between predicted and measured energy consumption.
@@
However, a number of approaches and tools capable of proactively estimating the energy and environmental cost of training have also emerged. LLMCarbon [13] is a very recent end-to-end CF projection model designed for both dense (all parameters are used for every input) and mixture of experts (only a subset of parameters for each input) LLMs. It incorporates critical LLM, hardware, and data center parameters, such as LLM parameter count, hardware type, system power, chip area, and data center efficiency, to model both operational and embodied CFs of a LLM. Furthermore, [14] aims at providing guidelines for scaling AI in a sustainable way. They analyze the energy efficiency of new processing units, the CF of the most prominent LLM models since GPT-3 and analyze their lifecycle carbon impact showing that inference and training are comparable. They conclude that, to enable sustainability as a computer system design principle, better tools for carbon telemetry are required, large scale carbon datasets, carbon impact disclosure and more suitable carbon metrics.
@@
The networking and IoT communities, increasingly relying on AI techniques, are also investigating pathways towards net-zero carbon emissions [15]. In [15] the authors notice that in spite of improvements in hardware and software energy efficiency, the overall energy consumption of mobile networks continues to rise, exacerbated by the growing use of resource-intensive AI algorithms. They introduce a novel evaluation framework to analyze the lifecycle of network AI implementations, identifying major emission sources. They propose the Dynamic Energy Trading and Task Allocation framework, designed to optimize carbon emissions by reallocating renewable energy sources and distributing tasks more efficiently across the network. Similar as in [14], they highlight the need for the development of new metrics to quantify the environmental impact of new network services enabled by AI.
@@
The authors of [16] introduce a novel framework to quantify energy consumption and carbon emissions for vanilla FL methods and consensus-based decentralized approaches, identifying optimal operational points for sustainable FL designs. Two case studies are analyzed within 5G industry verticals: continual learning and reinforcement learning scenarios. Similar to the authors of [15], they consider energy [$Wh$] and [$CO_{2}eq$] in their evaluations. The solution proposed in [17] is a hardware/software co-design that introduces modality gating (throttling) to adaptively manage sensing and computing tasks. This so-called adaptive modality gating (AMG) solution features a novel decoupled modality sensor architecture that supports partial throttling of sensors, significantly reducing energy consumption while maintaining data flow across multimodal data: text, speech, images, and video. More energy efficient task scheduling has been considered in  [18], while more efficient neural network architecture design and subsequent model development in [19]. The energy efficiency of various programming languages is studied in [30], while energy and carbon consumption testing for AI-driven IoT services is developed in [20].
@@
Furthermore, novel metrics to evaluate Deep Learning (DL) models, considering not only their accuracy and speed but also their energy consumption and cost have been proposed in [21]. The four metrics used are: Accuracy Per Consumption (APC) and Accuracy Per Energy Cost (APEC) for inference, and Time To Closest APC (TTCAPC) and Time To Closest APEC (TTCAPEC) for training. These metrics aim to promote the use of energy-efficient and green energy-powered DL systems by integrating energy considerations into the performance evaluation. Nevertheless, while powerful in terms of assessing the energy/performance trade-offs of models, these metrics are less suitable in assessing the energy efficiency of a system such as AIoT.
Inspired by the simplicity and effectiveness of metrics such as transmission efficiency through energy-per-bit and the lifecycle emissions, and responding to calls for new metrics [15, 14] that capture the energy cost in the era of AI enabled systems, this work proposes the novel Energy Cost of AIoT Lifecycle (eCAL) metric and methodology to derive it. eCAL is the only metric that, based on OSI and MLOps architectures, provides insights into the cost of inference considering the data transmission and AI model overheads.
III eCAL Definition and Methodology
In our work, we consider an AIoT communication and computing system whose lifecycle is depicted in Fig. 1, consisting of the following data manipulating components as per the OSI stack and MLOps architectures:
@@
This data-manipulating component, depicted in the lower part of Fig. 1 includes receiving telemetry data from sensors $N_{\mathrm{S}}I_{\mathrm{S}}$ to the terminal computing infrastructure via wired or wireless technologies. The collected information can be turned into indicators such as signal strength and be interpreted by the AI models, for example, to predict the location of a user. To successfully collect all the application level data, the total energy of data collection $E_{\mathrm{DC}}\leavevmode\nobreak\ [J]$ is needed.
Figure 1: Data manipulation components involved in the lifecycle of an AIoT system.
@@
To ensure accuracy and reliability during the training process, the data must go through several preprocessing steps such as cleaning, feature engineering and transformation. The energy consumption for preprocessing $E_{\mathrm{pre}}$, depends on the integrity of the ingested dataset.
@@
The training component comprises the AI/ Machine Learning (ML) model development using selected AI/ML techniques, such as neural network architectures, and data. In this step of the model development process, the processed data, $N_{\mathrm{S,T}}$ is fetched and utilized to learn weights and biases in order to approximate the underlying distribution. For most neural architectures, the learning processes rely on Graphics Processing Unit (GPU) and Tensor Processing Unit (TPU) performing complex tensor processing operations and consuming training energy, $E_{\mathrm{train}}$. Once the neural network architecture weights are learned using the data in view of minimizing a loss function, the model is considered ready for evaluation and deployment. Subsequently, the quality of the learned model is evaluated on the evaluation datasets, $N_{\mathrm{S,E}}$, a process that consumes $E_{\mathrm{eval}}$ energy. The data storage, preprocessing, training, and evaluation form the end-to-end training of the AI/ML model and cumulatively require $E_{\mathrm{D}}$ energy to complete.
@@
Once the model is trained, it can be used by applications in inference mode. Various applications can send samples of data $N_{\mathrm{I,P}}$ and receive model outputs in the form of forecasts for regression problems or discrete (categorical) labels for classification problems. The energy consumption $E_{\mathrm{inf}}$ of a single inference is relatively small, however for high volumes of requests it can become significant.
@@
$N_{\mathrm{IoT,cycle}}$ No. of processing cycles per bit (IoT device)
@@
$B_{\mathrm{T,L_{OSI}}}$ No. of requested bits by the application from the IoT device
@@
$P_{\mathrm{IoT,cycle}}$ Power consumption per processing cycle (IoT device)
@@
$E_{\mathrm{T}}$ EC of transmitting data from IoT device to the reception side
@@
$eCAL_{\mathrm{abs}}$ EC over the lifetime of a model in the AIoT system
$eCAL$ EC per bit over the lifetime of a model in the AIoT system
@@
Based on this lifecycle, we proposed eCAL as the ratio between all the energy consumed by all data manipulation components and all the manipulated application-level bits:
@@
In the upcoming sections, we employ the following methodology in order to finally derive the proposed metric. We follow the data flow, highlighted with blue in Fig. 1, across the above-mentioned components and derive the computational complexity and subsequently the energy consumption, highlighted with purple, in each component. For instance, in Section IV, we derive $E_{\mathrm{T}}$ for the data collection component. With this formalism in place, Section VIII is concerned with deriving eCAL.

### v3

# The Energy Cost of Artificial Intelligence Lifecycle in Communication Networks

@@ Abstract
Artificial Intelligence (AI) is being incorporated in several optimization, scheduling, orchestration as well as in native communication network functions. While this paradigm shift results in increased energy consumption, quantifying the end-to-end energy consumption of adding intelligence to such systems is particularly challenging. Conventional metrics focus on either communication, computation infrastructure, or model development. To address this, we propose a new metric, the Energy Cost of AI Lifecycle (eCAL) of one AI model in a system. eCAL captures the energy consumption throughout the development and deployment of an AI-model providing intelligence in a wireless communication network by analyzing the complexity of data collection and manipulation in individual components and deriving overall and per-bit energy consumption. We show that the better a model is and the more it is used, the more energy efficient an inference is. For a simple case study, eCAL for making $100$ inferences is $2.73$ times higher than for $1000$ inferences. Additionally, we have developed a modular and extendable open-source simulation tool to enable researchers, practitioners, and engineers to calculate the end-to-end energy cost with various configurations and across various systems, ensuring adaptability to diverse use cases.
@@
AI Model lifecycle, Energy Consumption, Carbon Footprint, Metric, Methodology
@@ I Introduction
Next-generation wireless networks aim to deliver unprecedented levels of connectivity, service diversity and flexibility. Artificial Intelligence (AI) is expected to become fundamental to 6G and future cellular networks [1], with traditional networking functions increasingly being supplanted by AI‑driven implementations to realize so‑called “AI‑native” capabilities  [2, 3]. These AI techniques aim to deliver intelligent, sustainable, and dynamically programmable services that enhance network adaptability and flexibility [1]. Recent works propose a pervasive multi-level native AI architecture, which incorporates knowledge graph (KG) into mobile network reducing data scale and computational costs of AI training by almost an order of magnitude [4], attempt to improve communication network efficiency by minimizing the transmitted over the air data through semantic communications [5], optimize end-to-end network slicing [6]. Furthermore, and AI/ Machine Learning (ML) workflow was defined  [7] as part of the Open Radio Access Network (O-RAN) Alliance that tackles the standardization of the network access segment.
While such systems designed through endogenous AI models enable unprecedented automation and agility as well as improved energy consumption for the performed data transmission, their reliance on AI implies additional energy consumption and increased carbon emissions [8]. These energy and carbon costs may be significant [9], especially in the cases where models with many parameters are employed, such as large language models. Projections indicate that by 2030, electricity consumption could rise to 100 TWh for data centers and 40 TWh for telecommunication networks [10].
@@
In light of climate challenges, the general AI research community’s efforts are intensifying to better estimate the costs of training [11] and using [12] AI, the Carbon Footprint (CF) of LLMs [13] and find ways to scale models efficiently [14]. Furthermore, the networking community is increasingly relying on AI techniques, are also exploring pathways towards net-zero carbon emissions [15]. They analyze the CF of different learning techniques [16] and data modalities and develop different approaches to optimize energy efficiency, including joint optimization of hardware-software co-design [17], improved scheduling [18], more efficient model design [19], and energy/carbon consumption testing [20].
@@
Depending on the scope of the study, various metrics are employed to understand energy efficiency and CF. As well noted in [15], the traditional ITU standardized Energy-per-Bit $[J/b]$ metric ’will no longer be able to reflect the environmental impact of the modern mobile services, especially network AI-enabled smart services’. For instance, [17] considers normalized energy as well as energy for buffering in $[J]$, electricity $[Wh]$ and $[CO_{2}eq]$ for the various phases involved in federated learning (FL) edge systems [15, 16], $[J]$ and GFLOPS of the neural architectures [19] and energy $[Wh]$ for the scheduled loads in [18]. However, these metrics are not specifically designed to capture the energy efficiency of a system or its parts. To this end, a metric somewhat similar to the Energy-per-Bit $[J/b]$ would be more suitable for measuring energy efficiency of an AI-based communication system, similarly as some other efficiency metrics that have been proposed in wireless in particular and engineering and economy in general [21]. For instance, spectral efficiency, measured in $[b/s/Hz]$ [22] provides means of calculating the amount of data bandwidth available in a given amount of spectrum, and lifecycle emissions for vehicles, measured in $[g/km]$ [23], enables computing the CF in grams per kilometer.
@@
Following these observations, we tackle a timely, unexplored and complex problem aimed to find a suitable energy efficiency metric that would be simple and general and would holistically quantify the environmental cost of inference carried out in future communication networks. This work is meant as the foundation on which to build increasingly accurate understanding of the cost of AI model embedding in networks. The contributions of this paper are as follows:
@@
  • We propose a new metric, the Energy Cost of AI Lifecycle (eCAL), measured in $[J/b]$, that captures the overall energy cost of generating an inference in a communication system. Unlike capturing the energy required for transmitting bits that is done by the Energy-per-Bit $[J/b]$ metric, or $[J]$ for AI model complexity, the proposed eCAL captures the energy consumed by all the data collection and manipulation components in an AI-enhanced system during the lifecycle of a trained model delivering inference.
@@
  • Following the Open Systems Interconnection (OSI) and AI/ML workflow or Machine Learning Operations (MLOps) [24] architectures, we devise a methodology for determining the eCAL of an communication system and develop an open source modular and extensible simulation tool [^1] https://github.com/sensorlab/eCAL, namely, eCAL calculator. The methodology breaks down the system into different data manipulation components, and, for each component, it analyzes the complexity of data manipulation and derives the overall energy consumption.
@@
  • Using the proposed metric, we demonstrate that the better a model is and the more it is used, the more energy-efficient an inference is. For a selected case study, the energy consumption per bit for making $100$ inferences is $2.73$ times higher than for $1000$ inferences. This can be explained by the fact that the energy cost of data collection, preprocessing, training, and evaluation, all necessary to develop an AI model are spread across more inferences.
@@
The paper is structured as follows. We first review the relevant literature in Section II. We then describe the lifecycle of an AI model, introduce eCAL and the methodology for deriving it in Section III. Subsequently, the derivation of the energy consumption formulas for different data manipulation components involved in the AI model lifecycle are provided in Sections IV to VII. In Section VIII, the proposed metric, eCAL, is derived and analyzed for selected communication and model development cases, demonstrating its utility in capturing the overall energy cost of an AI model over its lifecycle. Finally, Section IX concludes the paper and outlines future research directions.
@@
Proposed Metric eCAL $J/b$ Quantifies the energy consumption of AI inference, considering both communication and computation energy costs relative to application data.
@@ II Related Work
The growing dependence on AI has introduced significant energy and carbon costs [8], intensifying environmental concerns [13]. In response, the AI research community has placed greater emphasis on understanding and then formalizing methods to assess the environmental impact of the increasing AI adoption. Due to the complexity of state-of-the-art neural architectures, in many cases the energy cost of training is computed after the training is done, by measuring the performed computation through interfaces such as the performance application programming interface [29]. Furthermore, [11] showed that the energy footprint of inference has been more extensively studied than that of training. The study highlighted that the accuracy between the predicted and measured energy consumption is below 70%.
@@
Meanwhile, a number of approaches and tools capable of proactively estimating the energy and environmental cost of training have also emerged. LLMCarbon [13] is a very recent end-to-end CF projection model designed for both dense (all parameters are used for every input) and mixture of experts (only a subset of parameters for each input) LLMs. It incorporates critical LLM, hardware, and data center parameters, such as LLM parameter count, hardware type, system power, chip area, and data center efficiency, to model both operational and embodied CFs of a LLM. Furthermore, [14] aimed to provide guidelines for scaling AI in a sustainable way. They examined the energy efficiency of new processing units and the CF of the most prominent LLM models since GPT-3, analyzed their lifecycle carbon impact, and showed that inference and training are comparable. They concluded that to enable sustainability as a computer system design principle, better tools for carbon telemetry, large-scale carbon datasets, carbon impact disclosure, and more suitable carbon metrics are required.
@@
The networking community, increasingly relying on AI techniques, are also investigating pathways towards net-zero carbon emissions [15]. In [15] the authors noticed that in spite of improvements in hardware and software energy efficiency, the overall energy consumption of mobile networks continues to rise, exacerbated by the growing use of resource-intensive AI algorithms. They introduced a novel evaluation framework to analyze the lifecycle of network AI implementations, identifying major emission sources. They proposed the dynamic energy trading and task allocation framework, designed to optimize carbon emissions by reallocating renewable energy sources and distributing tasks more efficiently across the network. Similarly to [14], they highlighted the need for the development of new metrics to quantify the environmental impact of new network services enabled by AI.
@@
The authors of [16] introduced a novel framework to quantify energy consumption and carbon emissions for vanilla FL methods and consensus-based decentralized approaches, and identified optimal operational points for sustainable FL designs. Two case studies were analyzed within 5G industry verticals: continual learning and reinforcement learning scenarios. Similar to the authors of [15], they considered energy [$Wh$] and [$CO_{2}eq$] in their evaluations. The solution proposed in [17] is a hardware/software co-design that introduces modality gating (throttling) to adaptively manage sensing and computing tasks. This so-called adaptive modality gating (AMG) solution features a novel decoupled modality sensor architecture that supports partial throttling of sensors, significantly reducing energy consumption while maintaining data flow across multimodal data: text, speech, images, and video. More energy efficient task scheduling has been considered in  [18], while more efficient neural network architecture design and subsequent model development in [19]. The energy efficiency of various programming languages was studied in [30], while energy and carbon consumption testing for AI-driven Internet of Things (IoT) services were developed in [20].
@@
Furthermore, novel metrics to evaluate Deep Learning (DL) models, considering not only their accuracy and speed but also their energy consumption and cost have been proposed in [28]. The four metrics used are: Accuracy Per Consumption (APC) and Accuracy Per Energy Cost (APEC) for inference, and Time To Closest APC (TTCAPC) and Time To Closest APEC (TTCAPEC) for training. These metrics aim to promote the use of energy-efficient and green energy-powered DL systems by integrating energy considerations into the performance evaluation. Nevertheless, while powerful in terms of assessing the energy/performance trade-offs of models, these metrics are less suitable in assessing the energy costs of adding intelligence to future communication networks.
Inspired by the simplicity and effectiveness of metrics such as transmission efficiency through energy-per-bit and the lifecycle emissions, and responding to calls for new metrics [15, 14] that capture the energy cost in the era of AI enabled systems, this work proposes the novel Energy Cost of AI Lifecycle (eCAL) metric and methodology to derive it. eCAL is the only metric that, based on OSI and MLOps architectures, provides insights into the cost of inference considering the data transmission and AI model overheads.
III Definition and Methodology
In our work, we consider a communication and computing system that enables the development of an AI model, whose lifecycle is depicted in Fig. 1, consisting of the following data manipulating components as per the OSI stack and MLOps architectures:
@@ III Definition and Methodology
This data-manipulating component, depicted in the lower part of Fig. 1, includes receiving data (i.e. telemetry, applications, etc) from devices $N_{\mathrm{S}}I_{\mathrm{S}}$ to the target computing infrastructure via wired or wireless technologies. The collected information can be turned into indicators such as signal strength and be interpreted by the AI models, for example, to predict the location of a user. To successfully collect all the application-level data, the total energy of data collection $E_{\mathrm{DC}}\leavevmode\nobreak\ [J]$ is needed. The number, topology or other particularities of the data collection component of the AI model lifecycle can be adapted to specific configurations if needed.
Figure 1: Data manipulation components involved in the lifecycle of an AI model.
@@
To ensure accuracy and reliability during the training process, the data must go through several preprocessing steps, such as cleaning, feature engineering and transformation. The energy consumption for preprocessing, $E_{\mathrm{pre}}$, depends on the integrity of the ingested dataset. The metric is conceptually capable of support also distributed or hybrid pre-processing as well as training and evaluation scenarios.
@@
The training component comprises the AI/ML model development using selected AI/ML techniques, such as neural network architectures, and the training data. In this step of the model development process, the processed data, $N_{\mathrm{S,T}}$ is fetched and utilized to learn weights and biases in order to approximate the underlying distribution. For most neural architectures, the learning processes rely on Graphics Processing Units and Tensor Processing Units, which perform complex tensor processing operations and consuming training energy, $E_{\mathrm{train}}$. Once the neural network architecture weights are learned using the data in view of minimizing a loss function, the model is considered ready for evaluation and deployment. Subsequently, the quality of the learned model is evaluated on the evaluation dataset, $N_{\mathrm{S,E}}$, a process that consumes $E_{\mathrm{eval}}$ energy. The data storage, preprocessing, training, and evaluation form the end-to-end training of the AI/ML model and cumulatively require $E_{\mathrm{D}}$ energy to complete.
@@
Once the model is trained, it can be used by applications or within network scheduling and optimization components in inference mode. Samples of data can be sent to the AI model running in inference mode $N_{\mathrm{I,P}}$ and model outputs are returned as responses in the form of forecasts for regression problems or discrete (categorical) labels for classification problems. The energy consumption $E_{\mathrm{inf}}$ of a single inference is relatively small, however, for high volumes of requests, it can become significant.
@@
$N_{\mathrm{dev,cycle}}$ No. of processing cycles per bit (device)
@@
$B_{\mathrm{T,L_{OSI}}}$ No. of requested bits by the application from the device
@@
$P_{\mathrm{dev,cycle}}$ Power consumption per processing cycle (device)
@@
$E_{\mathrm{T}}$ EC of transmitting data from device to the reception side
@@
$eCAL_{\mathrm{abs}}$ EC over the lifetime of an AI model in the system
$eCAL$ EC per bit over the lifetime of an AI model in the system
@@
Based on this lifecycle, we propose eCAL as the ratio of total energy consumed by data manipulation components to the total manipulated application-level bits:
@@
In the upcoming sections, we employ the following methodology in order to derive the proposed metric for example case studies that comply with the OSI and MLOps architectures. We follow the data flow, highlighted with blue in Fig. 1, across the above-mentioned components and derive the computational complexity and subsequently the energy consumption, highlighted with purple, in each component. For instance, in Section IV, we derive $E_{\mathrm{T}}$ for the data collection component. With this formalism in place, Section VIII is concerned with deriving eCAL under specific configuration assumptions, without limiting the scope of the proposed metric.